40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 12 23:54:10 2011, kiwamu, Update, IOO, EOM resonant box installed EOM.pngbefore.pngafter.png
    Reply  Wed Nov 6 20:34:28 2019, Koji, Update, IOO, Power combiner loss (EOM resonant box installed) power_combiner.pdf
This is a draft message, edit and submit it to make it permanent  
Message ID: 15018     Entry time: Wed Nov 6 20:34:08 2019     In reply to: 5391
Author: Koji 
Type: Update 
Category: IOO 
Subject: EOM resonant box installed 

 

Quote:

[Mirko / Kiwamu]

 The resonant box has been installed together with a 3 dB attenuator.

The demodulation phase of the MC lock was readjusted and the MC is now happily locked.

 

(Background)

We needed more modulation depth on each modulation frequency and so for the reason we installed the resonant box to amplify the signal levels.

Since the resonant box isn't impedance matched well, the box creates some amount of the RF reflections (#5339).

In order to reduce somewhat of the RF reflection we decided to put a 3 dB attenuator in between the generation box and the resonant box.

 

(what we did)

 + attached the resonant box directly to the EOM input with a short SMA connector.

 + put stacked black plates underneath the resonant box to support the wight of the box and to relief the strain on the cable between the EOM and the box.

 + put a 3 dB attenuator just after the RF power combiner to reduce RF reflections.

 + readjusted the demodulation phase of the MC lock.

 

(Adjustment of MC demodulation phase)

 The demodulation phase was readjusted by adding more cable length in the local oscillator line.

After some iterations an additional cable length of about 30 cm was inserted to maximize the Q-phase signal.

So for the MC lock we are using the Q signal, which is the same as it had been before.

 

 Before the installation of the resonant box, the amplitude of the MC PDH signal was measured in the demodulation board's monitor pins.

The amplitude was about 500 mV in peak-peak (see the attached pictures of the I-Q projection in an oscilloscope). Then after the installation the amplitude decreased to 400 mV in peak-peak.

Therefore the amplitude of the PDH signal decreased by 20 %, which is not as bad as I expected since the previous measurement indicated 40 % reduction (#2586).

 

 

ELOG V3.1.3-